Find the volume of a spherical wedge of radius 81 cm and central angle 19o.


1
Expert's answer
2021-07-05T04:52:22-0400

volume of a wedge = "\frac{1}{270} \times R^3 \times\theta"


= "\frac{1}{270} \times 81^3 \times19"


= 37397.7 cm"^3"